I Review I ConstantsI Periodic Table PartA A child on a swing set swings back and forth with a period of 3.0 s and an amplitude of 25°. Suppose that the motion is simple harmonic. What is the maximum speed of the child as she swings? Express your answer with the appropriate units. 四j ? 0.785 Umax
Answer #1

Here ,

let the maximum speed of child as the child swings is v

let the length of the string is L

T = 2pi * sqrt(L/g)

3 = 2pi * sqrt(L/9.8)

solving for L

L = 2.23 m

Now, Using conservation of energy

m * 9.8 * 2.23 * (1 - cos(25)) = 0.50 * m * v^2

9.8 * 2.23 * (1 - cos(25)) = 0.50 * v^2

v = 2.02 m/s

the maximum speed of the child is 2.02 m/s